OpenStudy (anonymous):

What is the solution to the rational equation 4/3 + 2/x+5 = -6/3x+15 ?

OpenStudy (australopithecus):

combine the fractions on both sides of the equation and bring x into the numerator

OpenStudy (australopithecus):

remember you need a common denominator to add or subtract fractions

OpenStudy (australopithecus):

First write 4/3 + 2/x+5 and -6/3x+15 as a single fraction

OpenStudy (anonymous):

okayy

OpenStudy (australopithecus):

sorry I probably didn't articulate that well enough write 4/3 + 2/x+5 as a single fraction and -6/3x+15 as a single fraction keep them seperate

OpenStudy (anonymous):

so 6/ 3x+15 = -6/ 3x+15 ?

OpenStudy (anonymous):

what is the value of x

OpenStudy (australopithecus):

One side of the equation simplified 4/3 + 2/x+5 = 4(x+5)/3(x+5) + 3(2)/3(x+5) = ( 4(x+5) + 3(2) )/3(x+5) The otherside simplified -6/3x+15 = -6/3x + 15(3x)/3x = (-6 + 15(3x))/3x now you have (-6 + 15(3x))/3x = ( 4(x+5) + 3(2) )/3(x+5) now bring the variables into the numerator (-6 + 15(3x))3(x+5)(3x)/3x = ( 4(x+5) + 3(2) )(3x)(x+5)/(x+5) = (-6 + 15(3x))3(x+5) = (4(x+5) + 3(2) )(3x) Now expand the terms and simplify you should be able to do this

OpenStudy (anonymous):

4/3 + 2/(x+5) = -6/(3x+15) ?

OpenStudy (anonymous):

@chanh_chung yea

OpenStudy (anonymous):

[4*(x+5)+2*3+6]/(3x+5)=0 12x+72/(3x+5)=0 12x+72=0 x=-72/12=-6

OpenStudy (anonymous):

value of x A. -2 B. 2 C 8 D =8

OpenStudy (anonymous):

D. -8 *

OpenStudy (anonymous):

you got x = -72 ?

OpenStudy (anonymous):

sr

OpenStudy (anonymous):

what ? i gave the four answer choices

OpenStudy (anonymous):

sorry [4*(x+5)+2*3+6]/(3x+5)=0 (4x+32)/(3x+5)=0 4x+32=0 x=-32/4=-8

OpenStudy (anonymous):

okayy thank you